Calculate the osmotic pressure of a 5% solution of glucose (mol. wt=180) at 291 k

Grade:12

1 Answers

Haridharan Srinivasa Desigan
6 years ago
π=iMRTi=1(asglucose is a non electrolyte dissolved in water) ,M=[(w/W%)/100]*1 = 0.05mol/mL ,R=0.0826,T=281K>>> π = 1*0.05*0.0826*291>>> π = 1.202 >>> osmotic pressure of glucose solution is π = 1.202 !!